OpenStudy (anonymous):

Don't Understand?.... (Pre-Cal) Evaluate for Sin 2u, cos 2u, tan 2u... if Tan u= -4 ? How do I solve for Tan -4? what formula?

OpenStudy (anonymous):

am i supposed to start with sin -4/cos -4?

OpenStudy (anonymous):

do i have to use a half angle formula to find the result for sin or cos?

ganeshie8 (ganeshie8):

use below \[\sin(2u) = \frac{2\tan (u)}{1+\tan^2 (u)}\] \[\cos(2u) = \frac{1-\tan^2 (u)}{1+\tan^2 (u)}\]

OpenStudy (anonymous):

how did you derive the first formula?

OpenStudy (anonymous):

im still confused on how to find the values?

OpenStudy (anonymous):

nevermind i think i got it. thank you

ganeshie8 (ganeshie8):

np :) just plugin the tanu value and simlify